METCRAX: Logbook Entries

METCRAX: Site West_Lower Messages, 13 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
185 Tue 15-May-2007Horizontal Rnet at wl had wrong coefficientsWest_LowerSemmer
160 Wed 01-Nov-2006wl teardownWest_LowerOncley
129 Wed 25-Oct-2006krypton kleanWest_LowerOncley
109 Sat 21-Oct-2006wl.0.5m sonic upWest_LowerOncley
103 Fri 20-Oct-2006swapped wl.0.5m sonicWest_LowerOncley
67 Fri 13-Oct-2006wl sensor serial numbersWest_LowerMartin
52 Wed 11-Oct-2006TRH007 installed at wuWest_LowerHorst
45 Sun 08-Oct-2006data_stats bad at wl, restarted wlWest_LowerMartin
38 Sat 07-Oct-2006Installed sonic s/n 0673 at 0.5m, wlWest_LowerHorst
35 Fri 06-Oct-20060.5m sonic removed from wl.West_LowerMartin
34 Fri 06-Oct-20061.5m TRH reinstalled at wlWest_LowerMartin
33 Fri 06-Oct-2006gps fixed at wlWest_LowerMartin
7 Thu 05-Oct-2006Moved 1.5 m TRH to 0.5 m at wlWest_LowerHorst

185: Radiometers, Site West_Lower, Tue 15-May-2007 15:02:51 MDT, Horizontal Rnet at wl had wrong coefficients
I decided to use west lower as the radiation system for CHATS and did a check 
of the coefficients. I notice an error in the values used for the horizontal Q7.They had coefficients of 8.57 (+) and 10.53(-). The values should have been 
9 (+) and 11.22 (-). I am assuming that the Q7 locations were not changed 
during setup. 

160: LOG, Site West_Lower, Wed 01-Nov-2006 16:29:06 MST, wl teardown
West upper was torn down between ~1230-1500.

Serial numbers are:	16945	29136
PSP.out	16107
Pyg.out	27907
Tsurf	3577-3
Rnet parallel	91096
Rnet horizontal	Q94197
Gsoil	H983213
Tsoil	STP94017
TP01	200236
Logger "West Upper" 1829

0.5m	CSAT	367
1.5m	CSAT	366
3m	CSAT	1119
	Kh2o	1391
5m	CSAT	1117
8.5m	CSAT	674 [*** has loose hardware in electronics box!! *****]
Bridge	F

129: KH2O, Site West_Lower, Wed 25-Oct-2006 14:06:08 MST, krypton klean
krypton kleaned ~1145.  MSB of counts went from 03 -> 23.

I had to restart data system code (adn/aup) to get rserial to work.  (But it
seems like the archiving had been fine.)

109: Sonic, Site West_Lower, Sat 21-Oct-2006 15:44:10 MST, wl.0.5m sonic up
We got the 0.5m sonic going today.  With the Toughbook, it was possible to see
that the sonic was in fact just fine.  (Too bad we lugged the other PennState
spare all the way down!!)  We determined that the problem was with serial
port 5.  We switched to port 16 (which was somewhat difficult to reconfigure
using just the PDA!)  It turned out that the PDA cable didn't work for the
laptop, so we couldn't use the Toughbook to reconfigure.  Oh well...

103: Sonic, Site West_Lower, Fri 20-Oct-2006 18:23:39 MST, swapped wl.0.5m sonic
We swapped this sonic (637?) with 367 that just arrived from PennState, since
it had a serializer in it that we needed at rim.  However, we couldn't 
configure it and it now puts out a strange message with the status "nibble"
reporting a value of "E".  I also added a strange CSATB serial sensor on
the station's copy of isff.xml (not a full config) to use an end-message 
of 55ab, not 55aa.

Obviously this is a sick sonic.  The other PennState loaner (376) seemed to 
work much better in the trailer.

We did the swap about 11:00, but didn't get the new configuration working until
we were back in the base trailer, at about 15:30.


The sonic removed from wl was s/n 673; see logbook entry 38.

67: Station_Visit, Site West_Lower, Fri 13-Oct-2006 17:25:23 MST, wl sensor serial numbers
Q7		Q91096
Q7horiz		Q94197
pygin		EP29136
pygout		EP27907
pyrin		EP161945
pyrout		EP16107
Tsoil		STP94017
HFT		H983213
TP01		200236

52: TRH, Site West_Lower, Wed 11-Oct-2006 10:21:07 MST, TRH007 installed at wu
Replaced TRH007 at 1.5m on wu on the afternoon of 10/10/06.

Would not go back into the boom so used a file,
WD-40, and a small sledge to get it in... Might be 
still a little tight coming out.

45: Adam, Site West_Lower, Sun 08-Oct-2006 11:32:55 MST, data_stats bad at wl, restarted wl
The data_stats at wl indicate a problem, 
so the dsm software was stopped and restarted.

38: Sonic, Site West_Lower, Sat 07-Oct-2006 17:51:44 MST, Installed sonic s/n 0673 at 0.5m, wl
Installed sonic s/n 0673 (from rim) at 0.5 m wl

35: Sonic, Site West_Lower, Fri 06-Oct-2006 19:27:47 MST, 0.5m sonic removed from wl.
The flakey 0.5m sonic at wl was removed and brought back to the base.
S/N 0856

34: TRH, Site West_Lower, Fri 06-Oct-2006 19:26:53 MST, 1.5m TRH reinstalled at wl
The 1.5m TRH sensor was reinstalled at wl.


Reinstalled at 13:06 (Arizona) 10/6/06

33: Adam, Site West_Lower, Fri 06-Oct-2006 19:26:01 MST, gps fixed at wl
The gps at wl came back to life after I cycled power to it.
7: TRH, Site West_Lower, Thu 05-Oct-2006 09:14:05 MST, Moved 1.5 m TRH to 0.5 m at wl
10/4/06, morning

Charlie and I tried to install the TRH that had been sent to Boulder at 0.5 m 
on the west lower slope station, but it did not work.  
The T and RH data were not correct and the fan did not work.

Moved the 1.5 m TRH down to 0.5 m and took the bad TRH back to the base.


TRH 004 at 1.5 m was replaced by TRH 704 at 18:23 (UTC?)